PATGEN(1) PATGEN(1) NAME patgen−generatepatternsforTeXhyphenation SYNOPSIS patgendictionaryfilepatternfilelogfiletranslatefile DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thepatgenprogramreadsthedictionaryfileandproducesthepatternfile,writingatranscript ofitsactionsonthelogfile.Thedictionaryfileshouldcontainalistofhyphenatedwordsand previously-generatedpatterns. Thepatternsthusgeneratedcanbereadbyinitexforuseinhyphenatingwords.Fora(very) longexampleofpatgen’soutput,see$TEXMFMAIN/generic/hyphen/hyphen.tex,whichcontains thepatternsTEXusesforEnglish.Atsomesites,patternsforseveralotherlanguagesmaybe available,andthelocaltexprogramsmayhavethempreloaded;consultyourLocalGuideoryour systemadministratorfordetails. Allfilenamesmustbecomplete;noaddingofdefaultextensionsorpathsearchingisdone. FILES $TEXMFMAIN/generic/hyphen/hyphen.tex PatternsforEnglish. SEEALSO FrankLiang,Wordhy-phen-a-tionbycom-puter,STAN-CS-83-977,StanfordUniversityPh.D. thesis,1983. DonaldE.Knuth,TheTEXbook,Addison-Wesley,1986,ISBN0-201-13447-0,AppendixH. AUTHORS FrankLiangwrotethefirstversionofthisprogram.PeterBreitenlohnermadeasubstantialrevisionin1991forTEX3.ThefirstversionwaspublishedastheappendixtotheTEXwaretechnical report,availablefromtheTEXUsersGroup.HowardTrickeyoriginallyportedittoUnix. Web2C7.32x 27December1992 1